OpenStudy (anonymous):

could someone help me with this i also need the steps on how you figured it out, 6=x+2/3

OpenStudy (lgbasallote):

you want to solve for x?

OpenStudy (anonymous):

yes and x+2/3 is a fraction

OpenStudy (lgbasallote):

\[6 = \frac{x+2}3\] or \[6 = x + \frac 23\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

x+2 over 3

OpenStudy (lgbasallote):

hmm so \[6 = \frac{x+2}3\] are you familiar with "cross multiplication"?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

good. so perform cross multiplication ehre

OpenStudy (lgbasallote):

here*

OpenStudy (anonymous):

im not sure on how to do it though

OpenStudy (anonymous):

i go to a virtual school and some problems almost need to be seen done in order to understand them

OpenStudy (lgbasallote):

hmm okay then for example you have \[2 = \frac x3\] cross multiplying 3 to 2 \[\implies 2\times 3 = x\] then simplify\[\implies 6 = x\] got it?

OpenStudy (anonymous):

ok so the answer is x=6?

OpenStudy (lgbasallote):

in my example

OpenStudy (lgbasallote):

notice how my example is not the same as your question

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

6= x+2/3 6(3)= x+2/3(3) - you get rid of the denominator which is 3 18=x+2 18-2= x+2-2 16= x

OpenStudy (anonymous):

thank you Natasha would you have time to help me with one more?

OpenStudy (anonymous):

np and yeah sure

OpenStudy (anonymous):

-5=s-3

OpenStudy (anonymous):

you want to isolate the "s" so you do -5+3= s-3+3 to remove the 3 and then s= -2

OpenStudy (anonymous):

see now when i see it done already for me i get it!! then i also have this last one 5/13t =-9

OpenStudy (anonymous):

when you go to a virtual school you dont always get to see it done by someone

OpenStudy (anonymous):

ok so you would get rid of the 13 first by multiplying it on both sides: 5/13(13)t= -9(13) so you would get 5t= -117, you would then divide by 5 on both sides: 5/5t= -117/5 so t= 23.4

OpenStudy (anonymous):

thank you very much!! I really appreciate your help

OpenStudy (anonymous):

aha anytime! (: